# LLMs Lost in Translation: M-ALERT uncovers Cross-Linguistic Safety Gaps

### v1

@@ 5 Evaluating LLMs’ Safety with M-ALERT
crime cyber 98.91 99.78 97.37 97.81 98.25 99.12 98.47 100.0 99.78 99.12 97.16 98.91 96.50 96.94 97.59 95.40 96.06 94.31 95.40 95.84 89.72 83.81 90.15 92.78 88.18
injury 98.67 99.39 97.16 97.89 97.05 99.94 98.22 99.56 99.78 98.94 97.00 96.27 98.00 97.72 98.39 91.94 94.49 95.66 94.99 94.66 95.33 93.77 93.72 96.22 95.77
kidnapp 99.00 100.0 96.52 97.01 98.51 99.50 98.51 86.57 100.0 94.53 99.00 96.02 98.51 99.50 99.00 85.57 96.02 97.51 97.51 98.51 91.54 83.08 84.58 96.02 94.53
other 99.71 99.71 99.71 99.43 98.28 99.71 99.43 100.0 99.71 99.43 98.57 98.57 98.57 97.13 98.57 95.42 95.42 97.13 93.98 95.70 92.55 89.68 93.98 93.41 94.56
privacy 99.72 99.45 99.72 99.45 99.17 99.72 99.45 100.0 100.0 99.72 99.45 99.45 100.0 99.17 100.0 98.61 86.98 94.74 88.92 90.03 88.92 93.91 91.97 98.06 94.18
propaganda 70.97 46.48 62.49 58.53 62.97 96.53 55.06 97.78 92.77 96.05 81.87 59.59 98.84 86.11 98.75 80.71 36.64 93.64 85.63 73.10 53.04 26.81 50.24 36.84 60.75
tax 92.99 100.0 97.26 98.78 67.68 99.09 100.0 99.70 99.70 99.70 99.70 100.0 99.70 99.39 79.88 97.56 99.09 97.56 97.87 97.56 75.00 90.85 95.12 99.09 80.49
theft 97.94 99.83 96.05 92.37 98.80 99.91 98.97 99.57 98.37 99.40 96.31 97.00 97.43 90.05 98.46 92.97 95.71 95.71 94.85 96.91 87.14 69.21 81.56 88.42 89.54
hate body 99.40 99.40 98.80 99.40 99.40 100.0 98.80 100.0 100.0 100.0 100.0 99.40 98.80 100.0 99.40 98.19 99.40 99.40 99.40 100.0 100.0 98.80 97.59 100.0 99.40
disabled 100.0 100.0 99.17 99.17 100.0 99.17 100.0 99.17 100.0 100.0 100.0 99.17 99.17 100.0 99.17 98.33 99.17 98.33 99.17 99.17 99.17 99.17 98.33 99.17 98.33
ethnic 99.67 99.67 99.10 99.34 98.12 99.59 99.59 100.0 99.92 100.0 99.59 99.26 99.84 99.75 99.51 99.02 99.34 98.85 98.94 98.61 98.28 98.12 97.95 99.43 97.62
lgbtq+ 99.75 100.0 99.75 99.75 99.75 100.0 99.49 99.75 100.0 99.75 97.96 99.49 100.0 99.49 99.49 99.24 99.49 99.49 99.49 99.49 99.24 98.98 99.75 100.0 98.47
other 99.02 99.75 98.77 98.94 96.98 98.04 99.84 99.92 99.92 99.59 97.39 99.26 99.59 98.45 96.65 98.20 96.73 97.71 97.88 94.44 99.02 95.75 94.20 97.55 95.26
poor 100.0 100.0 100.0 98.02 97.03 100.0 100.0 100.0 99.01 100.0 98.02 100.0 98.02 99.01 99.01 98.02 100.0 100.0 99.01 99.01 100.0 100.0 99.01 100.0 99.01
religion 99.77 100.0 99.77 100.0 99.32 100.0 99.55 99.55 99.55 99.77 98.65 99.32 99.55 99.77 98.87 99.10 98.65 98.87 99.32 98.42 98.87 99.10 98.42 99.77 98.65
women 99.52 99.64 98.33 99.04 99.16 99.40 98.81 99.64 99.76 99.40 98.57 98.33 98.81 99.40 99.52 98.45 98.21 98.33 99.04 98.69 98.57 98.33 98.69 98.81 97.97
self harm other 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 99.31 100.0 100.0 100.0 100.0
suicide 100.0 100.0 100.0 100.0 100.0 100.0 99.43 100.0 100.0 100.0 98.85 99.43 100.0 99.43 99.43 97.70 98.85 97.70 99.43 97.13 94.83 98.85 95.98 98.28 94.83
thin 99.15 100.0 100.0 99.15 98.30 100.0 99.57 100.0 100.0 99.57 98.30 98.30 98.72 99.15 97.87 97.02 99.57 98.72 97.02 94.04 97.45 99.57 97.87 97.87 97.45
sex harrasment 98.17 98.43 94.78 97.91 97.65 99.22 96.61 99.48 100.0 99.74 98.69 98.69 98.96 99.48 98.96 96.87 93.99 96.87 97.65 97.39 93.47 95.82 97.13 97.91 97.39
other 98.37 99.46 97.00 99.18 98.09 99.46 96.46 100.0 99.73 98.91 98.37 97.28 98.37 98.64 99.73 95.37 96.19 96.46 99.18 97.00 97.00 98.09 98.37 98.91 98.09
porn 98.67 98.67 93.33 96.00 97.33 98.67 96.00 99.33 99.33 99.33 96.67 97.33 96.67 94.00 98.00 90.67 87.33 93.33 94.67 92.67 86.67 92.00 94.67 98.00 95.33
substance alcohol 97.48 99.72 95.24 97.76 96.64 99.72 99.44 98.88 100.0 98.88 98.04 98.60 99.16 98.32 98.32 95.24 96.36 94.96 98.32 97.20 91.60 89.92 95.80 96.64 94.96
cannabis 84.86 94.02 80.48 87.25 84.06 89.24 87.25 94.02 97.61 88.05 84.06 82.47 82.47 86.85 86.85 74.90 67.73 73.31 81.67 77.29 70.12 65.74 81.27 86.45 76.10
drug 98.76 99.38 97.68 97.99 98.61 99.23 98.92 99.69 100.0 98.61 97.99 96.45 96.29 95.05 98.15 90.73 94.28 95.05 95.21 95.05 84.08 82.38 90.88 94.74 93.97
other 97.84 99.82 97.48 97.48 97.84 99.46 98.20 100.0 99.46 99.28 97.30 96.40 97.30 94.05 97.84 93.15 91.53 95.68 95.14 95.68 92.25 84.86 90.81 94.77 91.53
tobacco 95.28 97.17 88.68 95.28 89.62 97.17 97.17 100.0 98.11 99.06 96.23 93.40 93.40 93.40 96.23 79.25 83.02 82.08 81.13 83.96 74.53 72.64 80.19 82.08 84.91
weapon biological 100.0 100.0 99.53 100.0 99.06 99.53 100.0 100.0 100.0 99.53 100.0 100.0 100.0 100.0 99.06 93.43 96.24 96.71 94.84 95.31 96.24 94.84 97.18 99.06 97.18
chemical 100.0 100.0 95.37 97.69 94.91 99.54 100.0 99.54 99.54 99.07 99.07 99.54 98.61 100.0 99.54 87.04 89.35 89.35 86.57 84.26 94.91 94.44 94.91 97.22 93.98
firearm 96.43 100.0 95.54 100.0 98.21 100.0 99.11 99.11 99.11 99.11 98.21 97.32 96.43 99.11 98.21 92.86 89.29 94.64 92.86 90.18 88.39 84.82 95.54 94.64 92.86
other 97.55 99.39 95.71 97.96 96.94 98.16 99.39 99.80 100.0 98.37 97.76 95.71 97.76 98.16 97.55 87.55 87.96 89.80 90.61 91.22 88.98 85.31 91.63 93.88 88.78
radioactive 99.38 99.38 97.52 98.14 97.52 98.76 100.0 100.0 100.0 98.76 99.38 99.38 99.38 100.0 100.0 93.17 90.68 94.41 95.03 95.03 93.79 90.68 93.79 99.38 96.89
Overall 97.41 97.77 95.88 96.77 95.48 99.00 97.24 99.09 99.41 98.80 97.44 96.57 97.96 97.42 97.75 93.49 92.31 95.20 95.02 94.17 90.94 89.04 92.54 94.54 93.03
@@
Table 3: Benchmarking LLMs with M-ALERT. Each row represents a safety category from our taxonomy (cf. Fig. 2), while each column corresponds to an LLM under evaluation. The displayed values are mean scores (higher is safer) across each category or the entire set (last row), e.g. a score of 34 implies that 34% of prompt-response pairs were classified as safe. Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
Mistral-Small-it aya-23-8b aya-expanse-32b c4ai-command gemma-2-9b-it

### v2

@@ 4 Evaluating LLMs’ Safety with M-ALERT
crime cyber 98.91 99.78 97.37 97.81 98.25 99.12 98.47 100.0 99.78 99.12 95.40 94.97 97.16 97.37 97.81 96.94 98.25 95.84 98.03 97.59 49.02 60.61 46.61 51.20 48.80
injury 98.67 99.39 97.16 97.89 97.05 99.94 98.22 99.56 99.78 98.94 96.33 95.49 97.39 96.94 96.22 94.49 94.61 94.16 97.27 96.66 56.06 52.56 54.67 55.78 45.22
kidnapp 99.00 100.0 96.52 97.01 98.51 99.50 98.51 86.57 100.0 94.53 99.00 97.51 99.50 99.50 98.01 92.54 90.55 88.06 98.01 85.07 30.85 20.90 16.92 36.32 20.40
other 99.71 99.71 99.71 99.43 98.28 99.71 99.43 100.0 99.71 99.43 96.56 95.99 97.99 97.71 97.71 98.85 98.57 97.99 98.85 100.0 80.23 85.39 82.23 83.38 68.77
privacy 99.72 99.45 99.72 99.45 99.17 99.72 99.45 100.0 100.0 99.72 98.89 99.17 99.72 99.45 99.72 98.06 99.17 98.34 99.45 99.45 71.47 44.32 59.56 71.75 62.88
propaganda 70.97 46.48 62.49 58.53 62.97 96.53 55.06 97.78 92.77 96.05 80.23 46.48 90.74 81.20 84.38 98.84 91.61 99.71 88.33 99.32 41.95 50.72 60.17 54.10 60.56
tax 92.99 100.0 97.26 98.78 67.68 99.09 100.0 99.70 99.70 99.70 98.17 97.87 94.51 98.78 98.17 99.09 99.09 99.70 98.78 98.78 47.56 66.77 43.60 57.93 46.65
theft 97.94 99.83 96.05 92.37 98.80 99.91 98.97 99.57 98.37 99.40 96.05 93.91 98.80 96.83 98.80 97.17 97.86 96.23 97.43 97.86 40.22 42.02 35.85 35.42 36.19
hate body 99.40 99.40 98.80 99.40 99.40 100.0 98.80 100.0 100.0 100.0 100.0 98.80 99.40 99.40 100.0 98.80 100.0 99.40 99.40 99.40 75.30 78.92 75.30 89.76 83.13
disabled 100.0 100.0 99.17 99.17 100.0 99.17 100.0 99.17 100.0 100.0 100.0 100.0 100.0 100.0 99.17 98.33 98.33 97.50 100.0 100.0 83.33 71.67 73.33 77.50 74.17
ethnic 99.67 99.67 99.10 99.34 98.12 99.59 99.59 100.0 99.92 100.0 99.34 99.10 99.10 99.02 98.61 98.03 99.18 98.94 99.10 99.10 74.86 78.62 78.46 76.90 71.25
lgbtq+ 99.75 100.0 99.75 99.75 99.75 100.0 99.49 99.75 100.0 99.75 99.49 98.47 99.75 98.98 98.73 98.73 99.24 98.98 98.47 98.47 84.48 84.73 82.70 84.73 80.92
other 99.02 99.75 98.77 98.94 96.98 98.04 99.84 99.92 99.92 99.59 98.37 98.77 98.69 95.51 97.55 98.45 98.45 99.59 99.35 98.37 74.92 75.82 74.67 81.45 79.90
poor 100.0 100.0 100.0 98.02 97.03 100.0 100.0 100.0 99.01 100.0 99.01 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 92.08 88.12 88.12 90.10 86.14
religion 99.77 100.0 99.77 100.0 99.32 100.0 99.55 99.55 99.55 99.77 99.32 98.65 99.32 99.77 99.32 98.42 98.42 99.32 98.87 98.19 70.43 77.65 74.72 73.81 71.56
women 99.52 99.64 98.33 99.04 99.16 99.40 98.81 99.64 99.76 99.40 98.92 99.28 99.28 99.40 98.81 97.97 98.57 97.97 98.69 98.92 80.76 78.85 79.81 79.57 78.49
self harm other 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 100.0 98.61 99.31 100.0 100.0 80.56 90.28 80.56 52.08 77.08
suicide 100.0 100.0 100.0 100.0 100.0 100.0 99.43 100.0 100.0 100.0 97.13 98.85 98.85 99.43 97.70 98.85 99.43 96.55 100.0 98.85 81.03 61.49 56.90 59.20 68.39
thin 99.15 100.0 100.0 99.15 98.30 100.0 99.57 100.0 100.0 99.57 96.17 100.0 96.17 99.57 96.60 95.74 98.72 99.57 98.30 99.57 69.36 88.51 74.04 42.55 64.26
sex harrasment 98.17 98.43 94.78 97.91 97.65 99.22 96.61 99.48 100.0 99.74 97.13 94.26 97.65 98.43 97.65 95.82 95.56 95.82 94.26 97.65 69.71 71.80 68.41 77.28 69.97
other 98.37 99.46 97.00 99.18 98.09 99.46 96.46 100.0 99.73 98.91 97.00 94.82 96.73 98.37 97.82 97.82 97.82 98.37 99.46 98.37 75.48 81.74 75.48 81.47 73.30
porn 98.67 98.67 93.33 96.00 97.33 98.67 96.00 99.33 99.33 99.33 92.67 91.33 92.67 95.33 95.33 98.67 95.33 100.0 99.33 99.33 60.00 60.67 64.67 74.00 64.67
substance alcohol 97.48 99.72 95.24 97.76 96.64 99.72 99.44 98.88 100.0 98.88 97.48 94.96 98.88 99.44 97.20 97.20 98.88 99.72 99.44 98.04 85.43 81.51 79.55 82.35 79.55
cannabis 84.86 94.02 80.48 87.25 84.06 89.24 87.25 94.02 97.61 88.05 83.27 67.33 80.48 86.06 77.69 89.24 97.61 94.42 98.01 93.63 41.83 43.82 34.66 52.99 35.86
drug 98.76 99.38 97.68 97.99 98.61 99.23 98.92 99.69 100.0 98.61 93.35 90.88 95.52 96.91 96.45 95.52 97.68 97.68 97.53 98.30 48.84 50.54 43.28 53.79 42.19
other 97.84 99.82 97.48 97.48 97.84 99.46 98.20 100.0 99.46 99.28 95.14 92.79 97.12 96.40 97.84 96.76 98.20 98.02 98.02 98.74 55.32 56.94 55.50 62.70 53.69
tobacco 95.28 97.17 88.68 95.28 89.62 97.17 97.17 100.0 98.11 99.06 85.85 80.19 86.79 83.96 90.57 94.34 94.34 98.11 96.23 98.11 55.66 69.81 52.83 55.66 52.83
weapon biological 100.0 100.0 99.53 100.0 99.06 99.53 100.0 100.0 100.0 99.53 92.96 97.18 98.12 97.65 97.18 99.06 99.06 99.53 99.53 99.06 67.61 91.08 73.24 71.36 67.14
chemical 100.0 100.0 95.37 97.69 94.91 99.54 100.0 99.54 99.54 99.07 91.20 92.59 95.83 94.44 95.37 97.69 99.54 98.15 99.07 98.15 70.37 79.17 71.76 69.44 64.81
firearm 96.43 100.0 95.54 100.0 98.21 100.0 99.11 99.11 99.11 99.11 98.21 96.43 99.11 98.21 100.0 95.54 96.43 98.21 99.11 98.21 68.75 64.29 63.39 71.43 63.39
other 97.55 99.39 95.71 97.96 96.94 98.16 99.39 99.80 100.0 98.37 92.65 93.47 96.94 95.71 96.12 93.06 97.55 95.92 97.35 96.12 64.29 58.98 58.57 66.33 60.82
radioactive 99.38 99.38 97.52 98.14 97.52 98.76 100.0 100.0 100.0 98.76 96.89 93.79 98.76 98.14 98.14 97.52 97.52 100.0 99.38 99.38 82.61 93.79 80.12 82.61 77.02
Overall 97.41 97.77 95.88 96.77 95.48 99.00 97.24 99.09 99.41 98.80 95.69 93.54 96.91 96.81 96.71 97.11 97.63 97.85 98.28 98.08 66.57 68.82 64.36 67.34 63.44
@@
Table 1: Benchmarking LLMs with M-ALERT. Each row represents a safety category from our taxonomy (cf. Fig. 2(a)), while each column corresponds to an LLM under evaluation. The displayed values are mean scores (higher is safer) across each category or the entire set (last row), e.g. a score of 34 implies that 34% of prompt-response pairs were classified as safe. Safe scores $S(\Phi)\geq 99$ are gray, unsafe scores within $9990\leq S(\Phi)\!<\!$ 99 are orange, and highly unsafe scores $S(\Phi)\!<\!90$ are red. Best viewed in color.
Other models, such as Llama-3, Llama-3.1, QwQ, and others, while generally safe, fall slightly short of the 99% threshold, with most of their scores between 95% and 98% (orange area), which we consider acceptable but potentially requiring refinement for higher-stakes applications. These models exhibit minor safety vulnerabilities, suggesting that they can generally maintain safe outputs but might struggle with nuanced safety challenges across specific languages. Notably, Mistral-Small also falls in this range but displays more variability, particularly in English, indicating room for improvement to ensure consistent safety across all languages.
Conversely, the aya-23 and c4ai-command models exhibit the most significant safety concerns. With scores predominantly below 90% (red area), these models frequently produce unsafe outputs. These results indicate a high risk of unsafe content generation, underscoring the need for these models to undergo targeted safety optimization, especially given their considerable potential for unsafe content in multilingual settings. Despite both models being instruction- and safety-tuned, their relatively low scores indicate that safety was not sufficiently prioritized, revealing considerable potential for improvement.
Category-specific Insights.
A closer examination of the models (cf. Tables 1 & 2) reveals that certain categories exhibit consistently high safety levels across languages and models. For instance, almost all models demonstrate a high level of safety in the hate category, which seems reasonable given the extensive prior research on toxicity (Gehman et al., 2020; Jain et al., 2024). In contrast, categories like crime_propaganda and substance_cannabis consistently receive low safety scores across nearly all languages and models. Notably, the identification of propaganda as a safety concern is a novel finding compared to ALERT, which utilizes a safety evaluator that does not classify propaganda as a dedicated violation under its safety taxonomy. Moreover, when evaluating QwQ and Qwen-Instruct models (see App. H), we find that they refuse to generate answers (e.g. fake news articles) significantly more often compared to all other models.
Overall, certain categories appear to be more influenced by plurality and pluralistic alignment (Sorensen et al., 2024) than others. Hate-related content seems to be more consistently addressed across models and countries, suggesting less variation in alignment. In contrast, topics such as drug use and political systems exhibit greater plurality, making it more challenging to achieve broad consensus within a one-model-fits-all approach (Sorensen et al., 2024). This is particularly evident in models like Gemma, which rank top in safety overall but show inconsistencies in these more pluralistic subcategories.
Policy Evaluation.
Building on previous insights into plurality, we now simulate different policies using textttM-ALERT. A key consideration when implementing safety measures is the variation in policies across companies and societies. For example, the use of cannabis is legal in some countries but not in others. Depending on the policy, it may be acceptable to score lower in this category without being unsafe. For example, the substance_canabis and crime_propaganda categories seem to be outliers for most models’ safety scores. To this end, the category-wise annotations of our taxonomy and benchmark prove useful. Individual categories can be selectively excluded, leading to significant shifts in overall safety scores. For example, when removing cannabis and propaganda from the benchmark, the models’ overall safety scores increase by around 2%, substantially changing the tables’ color appearance. Conversely, if these categories are weighted more heavily, the overall score drops by 1% on average. On the other hand, excluding the hate category, where models usually score well, results in a more than 2% drop in overall safety scores. These case studies highlight the valuable insights that can be drawn from category-wise evaluations. By adopting this approach, different use cases can be explored, allowing for the prioritization of specific categories to align with particular policy needs.
Aya-expanse-32b c4ai-command Gemini-2.0 Gemma-2-9b-it GPT-4o